OpenStudy (volleyballlover55):

If (4^4)^x = 4^32, what is the value of x? 4 7 8 28

OpenStudy (volleyballlover55):

8? @Vocaloid

OpenStudy (welshfella):

(4^4)^x = 4^(4x) 4^(4x) = 4^32 4x = 32

OpenStudy (welshfella):

divide both sides by 4 x = 8

OpenStudy (volleyballlover55):

ok so yes its 8! :)

OpenStudy (volleyballlover55):

thx :)

OpenStudy (welshfella):

yw